# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v14

@@ C.2 Geological Stratum
Sources: Created by the author. Data on the location of mines are obtained from the Association for Harmonious Cooperation (1932, 1937). Data on the geological stratum are obtained from the Ministry of Land, Infrastructure, Transport and Tourism.
@@
Sources: Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
@@ C.3 Dependent Variables
Sources: Data on the number of people and households are from the Statistics Bureau of the Cabinet (1935b, 1939). Data on the demographic variables are from the Statistics Bureau of the Cabinet (1932, 1938). Data on the labor force participation rates and employment shares are from the Statistics Bureau of the Cabinet (1933). Data on births, fetal, infant, and child deaths are from Aiikukai (1935); Social Welfare Bureau (1941).
@@ C.4.4 Elevation
Sources: See Sections C.4.1 to C.4.4 for the data on the distance and elevation variables. Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).
@@ D.1 Attenuation due to Measurement Error
Consider a simplified projection of $y_{i}$ on $\textit{MineDeposit}_{i}$ as $y_{i}=\theta\textit{MineDeposit}_{i}+\varepsilon_{i}$. When $\textit{MineDeposit}_{i}$ has a classical measurement error, $\textit{MineDeposit}^{*}_{i}=\textit{MineDeposit}_{i}+u_{i}$, where $u_{i}$ satisfies mean zero property. As explained in the main text, $u_{i}$ can be regarded as the unobservable changes in the assignments between the measured years of outcome and treatment variables. Given the random nature of the location of mines (Section 5.1), the measurement error ($u_{i}$) is likely to be uncorrelated with the outcome, true treatment variable, and error in the true model. Under this setting, the observable model becomes $y_{i}=\theta^{*}\textit{MineDeposit}^{*}_{i}+\nu_{i}$, where $\nu_{i}=\varepsilon_{i}-\theta u_{i}$. Consequently, I use $\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})=E[\textit{MineDeposit}^{*}_{i}\nu_{i}]=-\theta\text{var}(u_{i})\neq 0$ for $\theta^{*}=\theta+\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})/\text{var}(\textit{MineDeposit}^{*}_{i})$.
@@ D.2 Assumptions for the Identification
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+e_{i},\\ \text{{MineDeposit}}_{i}=\kappa+\tau\text{{Stratum}}_{i}+\epsilon_{i}.\end{split}}$ (20)
@@ D.3 Alternative Variance Estimator
D.4 Gender-Biased Structure Shift: Evidence from the 1930 Census
@@
D.5 Conditional Expectation of the Truncated Normal Distribution
@@
D.6 Testing the Validity of Threshold
@@
D.7 Sensitivity to the Regional Heterogeneity in Endowment of Male Miners
@@
D.8 Alternative Identification Strategy
@@
$\footnotesize{\begin{split}\textit{y}_{i,t}=\pi+\delta\textit{D}_{i,t}+\mathbf{x}^{\prime}_{i,t}\bm{\phi}+\eta_{i}+\lambda_{t}+\omega_{i,t},\end{split}}$ (21)
where $\eta$, $\lambda$, and $\omega$ are the municipal fixed effect, year fixed effect, and random error term, respectively. $\textit{D}_{i,t}$ is a treatment indicator variable that takes one for the treatment group in the post-treatment period and zero for otherwise. $\mathbf{x}$ includes the accessibility measures on the stations and sea ports. Accessibility measure on the rivers is not included as it is a time-constant variable. However, the municipal fixed effect effectively controls for all these time-constant unobservable factors, including the geological strata and terrain that were potentially correlated with the location of mines.
@@
D.9 Alternative Minerals: Heavy Metal Mines
@@
The validity of the main results on coal mines is assessed by comparing the results for heavy metal mines. Specifically, I focus on three heavy metals: gold, silver, and copper (GSC). This is because GSC mines comprise more than $80$% of all metal mines and generally extract these metals in one place. Although the AHC documented other types of metal mines, they were much smaller subsets: the number of lead and zinc, pig iron and steel, and tin mines recorded were only one, five, and four in 1931, respectively.
@@
D.10 Time-Series Plots of the IMR and FDR
@@
D.11 Heterogeneity with Respect to Scale of Mines
@@
  • (4) Association for Harmonious Cooperation. Zenkoku kōjyōkōzan meibo (The lists of factories and mines, 1932 edition). [in Japanese]. Association for Harmonious Cooperation, Tokyo, 1932.
@@
  • (6) Association for Harmonious Cooperation. Zenkoku kōjyōkōzan meibo (The lists of factories and mines, 1937 edition). [in Japanese]. Association for Harmonious Cooperation, Tokyo, 1937.
@@
  • (21)
  • (22) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, fukenhen (1930 Population Census of Japan, prefecture part). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1933
@@
  • (26) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, daiikkan (1930 Population Census of Japan, volume 1). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1935b.

### v15

@@ C.2 Geological Stratum
Sources: Created by the author. Data on the location of mines are obtained from the Cooperative Society (1932, 1937). Data on the geological stratum are obtained from the Ministry of Land, Infrastructure, Transport and Tourism.
@@
Sources: Data on the location of mines are from the Cooperative Society (1932, 1937).
@@ C.3 Dependent Variables
Sources: Data on the number of people and households are from the Statistics Bureau of the Cabinet (1935b, 1939). Data on the demographic variables are from the Statistics Bureau of the Cabinet (1932, 1938). Data on the labor force participation rates and employment shares are from the Statistics Bureau of the Cabinet (1935a). Data on births, fetal, infant, and child deaths are from Aiikukai (1935); Social Welfare Bureau (1941).
@@ C.4.4 Elevation
Sources: See Sections C.4.1 to C.4.4 for the data on the distance and elevation variables. Data on the location of mines are from the Cooperative Society (1932, 1937).
@@ D.1 Attenuation due to Measurement Error
Consider a simplified projection of $y_{i}$ on $\textit{MineDeposit}_{i}$ as $y_{i}=\vartheta\textit{MineDeposit}_{i}+\varepsilon_{i}$. When $\textit{MineDeposit}_{i}$ has a classical measurement error, $\textit{MineDeposit}^{*}_{i}=\textit{MineDeposit}_{i}+u_{i}$, where $u_{i}$ satisfies mean zero property. As explained in the main text, $u_{i}$ can be regarded as the unobservable changes in the assignments between the measured years of outcome and treatment variables. Given the random nature of the location of mines (Section 5.1), the measurement error ($u_{i}$) is likely to be uncorrelated with the outcome, true treatment variable, and error in the true model. Under this setting, the observable model becomes $y_{i}=\vartheta^{*}\textit{MineDeposit}^{*}_{i}+\nu_{i}$, where $\nu_{i}=\varepsilon_{i}-\vartheta u_{i}$. Consequently, I use $\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})=E[\textit{MineDeposit}^{*}_{i}\nu_{i}]=-\vartheta\text{var}(u_{i})\neq 0$ for $\vartheta^{*}=\vartheta+\text{cov}(\textit{MineDeposit}^{*}_{i},\nu_{i})/\text{var}(\textit{MineDeposit}^{*}_{i})$.
@@ D.2 Assumptions for the Identification
$\footnotesize{\begin{split}y_{i}=\alpha+\beta\text{{MineDeposit}}_{i}+e_{i},\\ \text{{MineDeposit}}_{i}=\iota+\zeta\text{{Stratum}}_{i}+\epsilon_{i}.\end{split}}$ (20)
@@ D.3 Alternative Variance Estimator
D.4 Evaluating the Effects of the Regulation on Mining Municipalities
As examined in Section 5.1, this study focuses on estimating the impacts of the coal mines on the local economy. The 1928 regulation included a grace period. Therefore, the data from the 1930 census and the 1933 vital statistics, which were collected before the regulation was enforced, feature municipalities with mines that either complied with or had not yet complied with the regulation. This means that when we consider an empirical setting for directly estimating the effects of the regulation, the “pretreatment” period cannot be precisely defined. In other words, the municipalities in the treatment group include some that were already treated during the “pretreatment” period. Therefore, the estimates from the standard difference-in-differences model are attenuated, because the time difference of the outcome variable in the treatment group cannot accurately capture the true increase or decrease in the outcome. Despite this attenuation issue, I use a policy evaluation model to examine whether fertility in the coal mining area increased after the regulation was enforced, because it relates to the main aim of this study.
First, I constructed a panel data set of units that were observed consistently in the analytical period. Thus, I excluded municipalities that were observed only in either the 1930 or 1935 census year. Next, I excluded the municipalities in the treatment group that were observed only in either the 1930 or 1935 census year. For the 1933 and 1938 vital statistics data, I applied the same trimming step. This process resulted in $1,094$ municipalities for both years ($2,188$ total observations).
Second, I consider the following structural equation for municipality $i$ in year $t$:
$\footnotesize{\begin{split}y_{i,t}=\varrho+\varpi\text{{MineDeposit}}_{i}\times\text{{Post}}_{t}+\mathbf{x}_{i,t}^{\prime}\bm{\varTheta}+\varsigma_{i}+o_{t}+\backepsilon_{i,t},\end{split}}$ (21)
where $y_{i,t}$ denotes the dependent variable, $\text{{MineDeposit}}_{i}$ indicates the indicator variable for mining area, $\text{{Post}}_{t}$ indicates the period after the enforcement of the regulation, $\varsigma_{i}$ indicates a municipal fixed effect, $o_{t}$ indicates a time fixed effect, identical to the post-treatment indicator, and ϶$\backepsilon_{i,t}$ is a random error term. Similar to the baseline cross-sectional analysis, I consider the reduced-form equation for the treatment variable ($\text{{MineDeposit}}_{i}$), as follows:
$\footnotesize{\begin{split}\text{{MineDeposit}}_{i}\times\text{{Post}}_{t}=\chi+\varkappa\text{{Stratum}}_{i}\times\text{{Post}}_{t}+\mathbf{x}_{i,t}^{\prime}\bm{\varPsi}+\varLambda_{i}+\varPhi_{t}+\eth_{i,t},\end{split}}$ (22)
where $\text{{Stratum}}_{i}$ indicates the stratum IV introduced in Section 5.1. $\varLambda_{i}$ and $\varPhi_{t}$ indicate a municipal- and a time-fixed effect, respectively, and ð$\eth_{i,t}$ is a random error term. The vector of control variables ($\mathbf{x}_{i,t}$) includes the distance to the nearest station and the distance to the closest port. The distance to the river and the average elevation considered in the cross-sectional analysis (Section 5.1) are not included, because they are time-constant variables.
Note that this model is designed to capture the effects of enforcing regulations on coal mining municipalities compared with the surrounding agricultural municipalities. It does not estimate the effects of the coal mine on the local economy. [^61] Online Appendix D.9 presents the exercise using a model that considers the opening of new coal mines as the treatment. Note too that it is not possible to trace the trends in the outcome variables in the pretreatment period, because the coal mining municipalities included in the treatment group differ from those in the previous census year of 1925. This is because some municipalities in the treatment group (measured in 1930 and 1935) had no mines in 1925. Additionally, there are no relevant vital statistics before 1933, meaning that the pretreatment trends for the early-life mortality variables are unobservable. Finally, because the municipalities included in the sample differ slightly from those used in the cross-sectional analysis, one must be careful when comparing these results with the baseline results presented in the main text.
Table D.2: Estimating the Effects of the Enforcement of the Regulation:
Evidence from the Panel Data Analysis
Panel A: Population, Fertility Responses, and Mortality Changes
Dependent Variable
[…]
$(0.036)$ $(0.860)$ $(8.884)$ $(1.522)$ $(1.439)$
Municipal and Year FEs Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes
Observations 2188 2188 2188 2188 2188
Clusters 1096 1096 1096 1096 1096
Years 1930 and 35 1930 and 35 1930 and 35 1930 and 35 1930 and 35
First-stage $F$-statistic 41.74 41.74 41.74 41.74 41.74
Panel B: Occupational Hazards and Social Changes
Dependent Variable
(1) IMR (2) FDR (3) Child MR (4) Sex Ratio (5) HH Size
[…]
$(13.517)$ $(8.451)$ $(3.386)$ $(0.023)$ $(0.071)$
Municipal and Year FEs Yes Yes Yes Yes Yes
Railway accessibility Yes Yes Yes Yes Yes
Port accessibility Yes Yes Yes Yes Yes
Observations 2188 2188 2188 2188 Yes
Clusters 1096 1096 1096 1096 Yes
Years 1933 and 38 1933 and 38 1933 and 38 1930 and 35 1930 and 35
First-stage $F$-statistic 41.74 41.74 41.74 41.74 41.74
Notes: This table shows the estimate ($\hat{\varpi}$) from the second-stage structural equation 21 in each outcome variable. The estimate ($\hat{\varkappa}$) from the first-stage reduced-form equation (22) is $0.325$ with the standard error of $0.050$. Panel A: Column 1 shows the result for the log-transformed population. Columns 2–5 show the results for the crude marriage rate, marital fertility rate, male mortality rate, and female mortality rate, respectively. Panel B: Columns 1-3 show the results for the infant mortality, fetal death, and child mortality rates, respectively. Columns 4–5 show the results for the sex ratio and average household size, respectively.
Table D.2 summarizes the results. In Panel A, column 1 presents an estimate for the population that is positive and weakly statistically significant, suggesting that coal mining municipalities experienced an approximate $7$ percentage point increase in population following the enforcement of the regulation. My baseline results show that the geometric mean of the treatment group increased from roughly $7,200$ to $7,700$ ($7$ percentage points), whereas the scale of the control group remained relatively stable (Table 2). The estimated magnitude from the panel data analysis aligns with the baseline results for the local population. Columns 2 and 3 reveal positive and statistically significant estimates for the crude marriage rate and marital fertility rate, respectively. These findings are in line with my baseline results in Table 3, confirming that the enforcement of the regulation led to an increase in fertility in the coal mining area.
The estimates for the crude mortality (columns 4–5 in Panel A), infant mortality (column 1 in Panel B), and fetal death (column 2 in Panel B) rates are all negative, but statistically insignificant. Overall, these findings are consistent with the baseline results, which show that mortality rates did not improve significantly in the coal mining area (Table 4). Although the baseline result suggests that early-life mortality moderately decreased in the coal mining area, this does not necessarily mean that all municipalities in the treatment group experienced a clear improvement in early-life deaths. [^62] Note that my baseline cross-sectional specification compares the sample means between the treatment and the control groups. In fact, the standard errors in columns 1 and 2 in Panel B are large relative to the estimated coefficients, which reflects heterogeneity in the variations among the treated units. The estimates for sex ratio and household size are negative, but statistically insignificant (columns 4 and 5 in Panel B), aligning with the main finding on social change that the influx of workers remained relatively stable.
D.5 Gender-Biased Structure Shift: Evidence from the 1930 Census
@@
D.6 Conditional Expectation of the Truncated Normal Distribution
@@
D.7 Testing the Validity of Threshold
@@
D.8 Sensitivity to the Regional Heterogeneity in Endowment of Male Miners
@@
D.9 Alternative Identification Strategy
@@
$\footnotesize{\begin{split}\textit{y}_{i,t}=\psi+\delta\textit{D}_{i,t}+\mathbf{x}^{\prime}_{i,t}\bm{\phi}+\eta_{i}+\lambda_{t}+\upsilon_{i,t},\end{split}}$ (23)
where $\eta$, $\lambda$, and $\upsilon$ are the municipal fixed effect, year fixed effect, and random error term, respectively. $\textit{D}_{i,t}$ is a treatment indicator variable that takes one for the treatment group in the post-treatment period and zero for otherwise. $\mathbf{x}$ includes the accessibility measures on the stations and sea ports. Accessibility measure on the rivers is not included as it is a time-constant variable. However, the municipal fixed effect effectively controls for all these time-constant unobservable factors, including the geological strata and terrain that were potentially correlated with the location of mines.
@@
D.10 Alternative Minerals: Heavy Metal Mines
@@
The validity of the main results on coal mines is assessed by comparing the results for heavy metal mines. Specifically, I focus on three heavy metals: gold, silver, and copper (GSC). This is because GSC mines comprise more than $80$% of all metal mines and generally extract these metals in one place. Although the CS documented other types of metal mines, they were much smaller subsets: the number of lead and zinc, pig iron and steel, and tin mines recorded were only one, five, and four in 1931, respectively.
@@
D.11 Time-Series Plots of the IMR and FDR
@@
D.12 Heterogeneity with Respect to Scale of Mines
@@
  • (40) Karl Pearson. On the criterion that a given system of deviations from the probable in the case of a correlated system of variables can be reasonably supported to have arisen from random sampling. Philosophical Magazine, 50(302):157-175, 1900.
  • (41)
@@
  • (4) Chikuhō Coal Mining Association. Chikuhō Sekitan Kōgyōkai Gojyūnen Shi (Fifty-Year History of Chikuhō Coal Mining Association). [in Japanese]. Chikuho Coal Mining Association, Fukuoka, 1935.
@@
  • (6) Cooperative Society. Zenkoku kōjyōkōzan meibo (The lists of factories and mines, 1932 edition). [in Japanese]. Cooperative Society, Tokyo, 1932.
  • (7)
  • (8) Cooperative Society. Zenkoku kōjyōkōzan meibo (The lists of factories and mines, 1937 edition). [in Japanese]. Cooperative Society, Tokyo, 1937.
@@
  • (16) Kazuichi Mikawa. Saishin saitan gaku, jyōkan (Latest Coal Mining Study, volume 1). [in Japanese] Shouhaku Shoin, Tokyo, 1939.
  • (17)
@@
  • (24) Statistics Bureau of the Cabinet. Taishōkyunen Kokuseichōsahōkoku, zenkokunobu dainikan (1920 Population Census of Japan, volume 2). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1929.
  • (25)
@@
  • (30) Statistics Bureau of the Cabinet. Shōwagonen Kokuseichōsahōkoku, dainikan (1930 Population Census of Japan, volume 2). [in Japanese]. Statistics Bureau of the Cabinet, Tokyo, 1935b.